Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Data collection)
mours, masserot - 12:39 Tuesday 14 May 2019 (45887) Print this report
Fixing the nanosecond field in a low latency h(t) file

The nanosecond field of the frame header was set to 100 instead of zero in the V1Online.ffl stream at GPS=1241800712 and 1241800751. This seems to be due to one the DAQ process (FsbAlp), being on an unusual state at that time.

The /data/prod/hrec/V1Online/V-V1Online-1241800000-2000.gwf file has been fixed today at 12:28 LT with frames having all nsec fields set to zero.

Comments to this report:
mours - 12:30 Wednesday 15 May 2019 (45897) Print this report

The /data/prod/hrec/V1Online/V-V1Online-1241800000-2000.gwf file was updated today once more to fix (set to zero) the timeOffset fields of the FrProcData of the two frames with problem.

Search
×

Warning

×